# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 6.5. Torus action and extension over the bogus cones
This follows immediately from Lemma 6.7– see also [48, Construction 1.24] in the general context of the intrinsic mirror construction. ∎
The action of the torus $T_{0}$ on the affine toric varieties $S_{\mathcal{X}^{\prime}}$ corresponding to the projective crepant resolutions $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ naturally induces an action of $T_{0}$ on the toric varieties $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ and $S^{\mathrm{sec}}_{\mathcal{X}}$, and on the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. It follows from the Equation (6.4) that the actions of $T_{0}$ on the families $\mathcal{Y}_{\mathcal{X}^{\prime}}\to S_{\mathcal{X}^{\prime}}$ given by Lemma 6.27 glue into an action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$.
@@
There exists a unique $T_{0}$ equivariant extension of the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$.
@@
This follows by the existence of the action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\to S_{\mathcal{X}}^{\mathrm{\mathrm{msec}}}$ and [50, Corollary 5.7]. ∎
Remark 6.29.
By Theorem 6.28, the line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ extends to a line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}}$. Moreover, by Lemma 6.27, the theta functions $\vartheta_{p}$ with $p\in P_{\mathbb{Z}}$ extend to sections of $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$. By Definition 6.26, every cone of the secondary fan $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$ is of the form $\rho+\sum_{v\in I}[X^{v}]$, where $\rho$ is a cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ and $I$ is a set of integral points $v\in P_{\mathbb{Z}}$ such that the corresponding irreducible components $X^{v}$ of $\mathcal{X}_{0}$ are either 1-surfaces or 2-surfaces. By [50, §6.4], in restriction to the toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ corresponding to the cone $\rho+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$ of $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, the divisor $\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}}$ given by Theorem 6.28 is given by the equation $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$.
@@ 7. KSBA stability of the mirror family to an open Kulikov degeneration
In this section we first prove the finite generation of their mirror algebras to open Kulikov degenerations, in Theorem 7.4. Then, in Theorem 7.7 we prove that the restriction of the mirror family to the dense torus in the base is KSBA stable. Finally, in Theorem 7.13 we prove the KSBA stability over the full base.
7.1. Finite generation
We show in Lemma 7.2 that the polarized mirror family to an open Kulikov degeneration restricts to a Mumford toric degeneration over a particular stratum of the base. This will be used in the following sections to prove finite generation.
7.1.1. Central fibers of polarized mirror families
Lemma 7.1.
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. The central fiber $(Y_{0_{\mathcal{X}}},L_{0,\mathcal{X}})$ of the polarized mirror family $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ over the unique torus fixed point $0_{\mathcal{X}}$ of the affine toric variety $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$ is the union of toric surfaces with intersection complex $(P,\mathscr{P})$. In particular, the set of irreducible components of $Y_{0_{\mathcal{X}}}$ is in one-to-one correspondence with $P_{\mathbb{Z}}$, all the irreducible components of $Y_{0_{\mathcal{X}}}$ are isomorphic to $\mathbb{P}^{2}$, and the line bundle $L_{0,\mathcal{X}}$ restricted to each irreducible component of $Y_{0_{\mathcal{X}}}$ is isomorphic to $\mathcal{O}_{\mathbb{P}^{2}}(1)$.
Proof.
This follows from [45, §2.1], where the central fibers of mirror families given by the intrinsic mirror construction are described explicitly. ∎
Lemma 7.2.
[…]
$\mathrm{Spec}\,\mathbb{C}[G]\cong\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})/J]\subset S_{\mathcal{X}}$
is a Mumford toric degeneration with central fiber $Y_{0_{\mathcal{X}}}$.
Proof.
As shown in §6.3, the wall-crossing transformations induced by the walls of the canonical scattering diagram are trivial modulo $J$, except for the contributions of the kinks corresponding to the finitely double curves $C$ of $\mathcal{X}_{0}$ with $F\cdot C=0$. The results follows by [41, §2.1] – see also [39, §6.2.1] and [16, §2]. ∎
7.1.2. Proof of finite generation
For every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the affine toric variety $\mathrm{Spec}\,\mathbb{C}[H]$ is a toric stratum of the affine toric variety $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$. We denote by $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ the corresponding open dense torus in $\mathrm{Spec}\,\mathbb{C}[H]$.
Lemma 7.3.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, for every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}\rightarrow T_{H}$ of the mirror family to $T_{H}$ is a finite type morphism, that is, $R_{\mathcal{X}}\otimes\mathbb{C}[H^{\mathrm{gp}}]$ is a finitely generated $\mathbb{C}[H^{\mathrm{gp}}]$-algebra.
Proof.
As $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a relative Mori dream space by Proposition 3.36, for every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$. By Equation (6.4), the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}\rightarrow T_{H}$ of the mirror family to $T_{H}$ is independent of the choice of the particular choice of crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$. By Lemma 6.12, there exists a crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that there exists a good divisor $F$ on $\mathcal{X}$ over $\mathcal{X}_{H}$. To prove Lemma 7.3, we will show that for such a particular crepant resolution, $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ is a finitely generated $\mathbb{C}[H]$-algebra, which implies in particular that $R_{\mathcal{X}}\otimes\mathbb{C}[H^{\mathrm{gp}}]$ is a finitely generated $\mathbb{C}[H^{\mathrm{gp}}]$-algebra.
Let $G$ be the face of $H$ spanned by the finitely many curves $C$ such that $F\cdot C=0$, and denote by $J:=H\setminus G$ the corresponding monoid ideal. By Lemma 7.2, $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ is a Stanley-Reisner algebra over $\mathbb{C}[H]/J\simeq\mathbb{C}[G]$, and so in particular a finitely generated $\mathbb{C}[H]/J$-algebra. In particular, there exists a finite set $\{\vartheta_{p}\}_{p\in I}$ of theta functions generating $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ as $\mathbb{C}[H]/J$-algebra. We will show that $\{\vartheta_{p}\}_{p\in I}$ actually generate $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ as $\mathbb{C}[H]$-algebra. Denoting by $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ the $\mathbb{C}[H]$-subalgebra of $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ generated by $\{\vartheta_{p}\}_{p\in I}$, we have to show that $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}=\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$. To do this, we follow the path of the proof of [41, Proposition 6.6] (see also [64, Lemma 16.6] and [71, Proposition 4.19]).
By Lemma 6.27, the divisor $-F$ induces an action of $\mathbb{C}^{\star}$ on $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ such that the monomials $t^{\beta}\vartheta_{p}$ with $\beta\in H$ and $p\in C(P)_{\mathbb{Z}}$ are eigenvectors with weight $-\varphi_{F}(p)-F\cdot\beta\in\mathbb{Z}_{\geq 0}$. For every $w,h\in\mathbb{Z}_{\geq 0}$, denote by $V_{w,h}$ the set of monomials $t^{\beta}\vartheta_{p}$ of weight $w$ with $p$ of height $h$, that is, with $p\in(hP)_{\mathbb{Z}}$. As the monomials $t^{\beta}\vartheta_{p}$ are a linear basis of $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$, it is enough to show that, for every $w,h\in\mathbb{Z}_{\geq 0}$, we have $V_{w,h}\subset(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$.
Fix $w,h\in\mathbb{Z}_{\geq 0}$. We prove that every monomial $t^{\beta}\vartheta_{p}$ in $V_{v,w}$ is actually in $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ by decreasing induction on $-F\cdot\beta$. First note that, as $(hP)_{\mathbb{Z}}$ is finite, there are finitely many possible values for $-\varphi_{F}(p)$, and so, as $w=-\varphi_{F}(p)-F\cdot\beta$ is fixed, there are finitely many possible values for $-F\cdot\beta$ for $t^{\beta}\vartheta_{p}$ in $V_{v,w}$. Hence, the result is immediate for large enough values of $-F\cdot\beta$. For the induction step, as $\{\vartheta_{q}\}_{q\in C(P)_{\mathbb{Z}}}$ generates $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$, one can write $\vartheta_{p}=a+b$, with $a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ and $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, and with both $a$ and $b$ homogeneous with respect to the $\mathbb{C}^{\star}$-action and of height $h$. Hence, we have $t^{\beta}\vartheta_{p}=t^{\beta}a+t^{\beta}b$, with $t^{\beta}a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and $t^{\beta}b\in V_{w,h}$. Actually, as $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, for every monomial $t^{\beta^{\prime}}\vartheta_{q}$ in $b$, we have $-F\cdot\beta^{\prime}\geq 1$. Hence, every monomial in $t^{\beta}b$ is of the form $t^{\beta+\beta^{\prime}}\vartheta_{q}$ with $-F\cdot(\beta+\beta^{\prime})>-F\cdot\beta$. It follows from the induction hypothesis that $t^{\beta}b\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and so finally that $t^{\beta}\vartheta_{p}\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$. ∎
Theorem 7.4.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the mirror algebra $\mathcal{R}_{\mathcal{X}}$ is a finitely generated $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$-algebra.
Proof.
As there are finitely many faces of $NE(\mathcal{X}/\overline{\mathcal{X}})$, it follows from Lemma 7.3 that there exists a finite set $\{\vartheta_{p}\}_{p\in I}$ such that, for every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, $\{\vartheta_{p}\}_{p\in I}$ generate $R_{\mathcal{X}}\otimes\mathbb{C}[H^{\mathrm{gp}}]$ as a $\mathbb{C}[H^{\mathrm{gp}}]$-algebra. Denote by $\mathcal{R}_{\mathcal{X}}^{\prime}$ the subalgebra of $\mathcal{R}_{\mathcal{X}}$ generated by $\{\vartheta_{p}\}_{p\in I}$. Both $\mathcal{R}_{\mathcal{X}}$ and $\mathcal{R}_{\mathcal{X}}^{\prime}$ are graded by the height function $h:C(P)_{\mathbb{Z}}\rightarrow\mathbb{Z}_{\geq 0}$. Hence, denoting for every $h\in\mathbb{Z}_{\geq 0}$ by $\mathcal{R}_{\mathcal{X},h}$ and $\mathcal{R}_{\mathcal{X},h}^{\prime}$ the subspaces of $\mathcal{R}_{\mathcal{X}}$ and $\mathcal{R}_{\mathcal{X}}^{\prime}$ of degree $h$, we have $\mathcal{R}_{\mathcal{X},h}^{\prime}\subset\mathcal{R}_{\mathcal{X},h}$. As $\mathcal{R}_{\mathcal{X},h}$ is a free finite rank $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$-module, with basis $\{\vartheta_{p}\}_{p\in(hP)_{\mathbb{Z}}}$, it follows that the quotient $\mathcal{R}_{\mathcal{X},h}/\mathcal{R}_{\mathcal{X},h}^{\prime}$ is a finitely generated $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$-module. On the other hand, for every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, we have $\mathcal{R}_{\mathcal{X},h}/\mathcal{R}_{\mathcal{X},h}^{\prime}\otimes\mathbb{C}[H^{\mathrm{gp}}]=0$. As the tori $\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ cover $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$, it follows from Nakayama’s lemma that $\mathcal{R}_{\mathcal{X},h}/\mathcal{R}_{\mathcal{X},h}^{\prime}=0$. Hence, $\mathcal{R}_{\mathcal{X},h}=\mathcal{R}_{\mathcal{X},h}^{\prime}$ for all $h$, and so $\mathcal{R}_{\mathcal{X}}=\mathcal{R}_{\mathcal{X}}^{\prime}$ is finitely generated. ∎
@@ 7.2. Generic KSBA stability of the mirror family
Proof.
First note that $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\rightarrow S_{\mathcal{X}}$ is flat since the algebra $\mathcal{R}_{\mathcal{X}}$ of theta functions is a free $NE(\mathcal{X}/\overline{\mathcal{X}})$-module. On the other hand, by Theorem 7.4, the algebra $\mathcal{R}_{\mathcal{X}}$ is finitely generated, and so the morphism $\mathcal{Y}_{\mathcal{X}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is projective.
To show that the fiber $(Y_{t},D_{t})$ over $t\in S_{\mathcal{X}}$ is semi-log-canonical and $K_{Y_{t}}+D_{t}=0$, consider the smallest toric stratum $\mathrm{Spec}\,\mathbb{C}[H]$ of $S_{\mathcal{X}}$ containing $t$, corresponding to a face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$. In particular, $t$ is contained in the dense torus orbit $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ in $\mathrm{Spec}\,\mathbb{C}[H]$. As in the proof of Lemma 7.3, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$, and the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}$ is independent of the choice of the crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$. Hence, by Lemma 6.12, we can assume that there exists a good divisor $F$ on $\mathcal{X}$ over $\mathcal{X}_{H}$. Let $G$ be the face of $H$ spanned by the curves $C$ with $F\cdot C=0$, and denote $J:=H\setminus G$. Consider the one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$ associated to $F$ as in §6.5. Consider the corresponding orbit $\mathbb{C}^{\star}\in T_{H}$ passing through $t$. Its closure in $\mathrm{Spec}\,\mathbb{C}[H]$ is an affine line $\mathbb{A}^{1}$ with $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 6.27, the restriction of $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})$ to this line is a $\mathbb{C}^{\star}$-equivariant family. Hence, by the proof of [43, Lemma 8.33], it is enough to show Theorem 7.5 for the central fiber of this family over $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 7.2, this central fiber is a fiber of a Mumford toric degeneration. Hence, the result follows from [6]. ∎
Theorem 7.6.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=0$.
Proof.
For the fibers of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, the result follows from Theorem 7.5 applied to the various projective crepant resolutions of $\overline{\mathcal{X}}$. By Theorem 6.28, any fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is related by $T_{0}$-action, and so is isomorphic, to a fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, and so the general result follows. ∎
Theorem 7.7.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every $t\in\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, let $(Y_{t},D_{t}+\epsilon\,C_{t})$ be the fiber over $t$ of the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. Then, the following holds:
@@
    If $t$ is contained in the dense torus orbit $T_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[N_{1}(\mathcal{X}/\overline{\mathcal{X}})]\subset\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, then the surface $Y_{t}$ is normal, the pair $(Y_{t},D_{t})$ is log canonical, and the pair $(Y_{t},D_{t}+\epsilon\,C_{t})$ is KSBA stable.
@@
We first prove i). By Theorem 7.6, the pair $(Y_{t},D_{t})$ is semi-log-canonical, and so log canonical since the surface $Y_{t}$ is assumed to be normal. Hence, by [50, Proposition 7.2], to prove that $(Y_{t},D_{t}+\epsilon\,C_{t})$ is semi-log-canonical, it is enough to show that $C_{t}$ does not contain any 0-dimensional stratum of $D_{t}$. By Remark 6.29, the divisor $C_{t}$ is defined by an equation of the form $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$ for some subset $I\subset P_{\mathbb{Z}}$ consisting of integral points corresponding to either 1-surfaces or 2-surfaces in $\mathcal{X}_{0}$. By Equations (6.8)-(6.9), we have $\vartheta_{p}|_{D_{t}}=0$ if $p\notin\partial P_{\mathbb{Z}}$. On the other hand, by Lemma 6.24, $D_{t}$ is a cycle of rational curves, with 00-dimensional strata $x_{v}$ in one-to-one correspondence with the integral points $v$ of $\partial P$ corresponding to the 0-surfaces in $\mathcal{X}_{0}$. Moreover, the theta functions $\vartheta_{p}$ with $p\in\partial P$ restrict to toric monomials on the irreducible components of $D_{t}$. It follows that for every 00-dimensional stratum $x_{v}$ of $D_{t}$, we have $\vartheta_{p}(x_{v})\neq 0$ if $p=v$, and $\vartheta_{p}(x_{v})=0$ if $p\in P_{\mathbb{Z}}\setminus\{v\}$. Hence, we have $\sum_{p\in P_{v}}\vartheta_{p}(x_{v})\neq 0$, and so $x_{v}\notin C_{t}$. This concludes the proof of i).

### v3

@@ 6.6. Torus action and extension over the bogus cones
This follows by the existence of the action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\to S_{\mathcal{X}}^{\mathrm{\mathrm{msec}}}$ and [53, Corollary 5.7]. ∎
Remark 6.42.
By Theorem 6.41, the line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ extends to a line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}}$. Moreover, by Lemma 6.26, the theta functions $\vartheta_{p}$ with $p\in P_{\mathbb{Z}}$ extend to sections of $\mathcal{L}_{\mathcal{X}}^{\mathrm{sec}}$. By Definition 6.40, every cone of the secondary fan $\mathrm{Mov}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ is of the form $\rho+\sum_{v\in I}[X^{v}]$, where $\rho$ is a cone of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ and $I$ is a set of integral points $v\in P_{\mathbb{Z}}$ such that the corresponding irreducible components $X^{v}$ of $\mathcal{X}_{0}$ are either 1-surfaces or 2-surfaces. By [53, §6.4], in restriction to the toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ corresponding to the cone $\rho+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$ of $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, the divisor $\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}}$ given by Theorem 6.41 is given by the equation $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. Then, the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=0$.
@@
For the fibers of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, the result follows from Theorem 6.36 applied to the various projective crepant resolutions of $\mathcal{X}^{\mathrm{can}}$. By Theorem 6.41, any fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is related by $T_{0}$-action, and so is isomorphic, to a fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, and so the general result follows. ∎
@@ 7. KSBA stability of the mirror family to an open Kulikov degeneration
In this section we first prove in Theorem 7.1 that the restriction of the mirror family to the dense torus in the base is KSBA stable. In Theorem 7.7 we extend this result to show KSBA stability over the entire base. Finally, in Theorem 7.8, we establish a result concerning the non-constancy of the mirror family.
@@ 7.1. Generic KSBA stability of the mirror family
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. For every $t\in\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, let $(Y_{t},D_{t}+\epsilon\,C_{t})$ be the fiber over $t$ of the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. Then, the following holds:
@@
    If $t$ is contained in the dense torus orbit $T_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]\subset\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, then the surface $Y_{t}$ is normal, the pair $(Y_{t},D_{t})$ is log canonical, and the pair $(Y_{t},D_{t}+\epsilon\,C_{t})$ is KSBA stable.
@@
We first prove i). By Theorem 6.43, the pair $(Y_{t},D_{t})$ is semi-log-canonical, and so log canonical since the surface $Y_{t}$ is assumed to be normal. Hence, by [53, Proposition 7.2], to prove that $(Y_{t},D_{t}+\epsilon\,C_{t})$ is semi-log-canonical, it suffices to show that $C_{t}$ does not contain any 0-dimensional stratum of $D_{t}$. By Remark 6.42, the divisor $C_{t}$ is defined by an equation of the form $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$ for some subset $I\subset P_{\mathbb{Z}}$ consisting of integral points corresponding to either 1-surfaces or 2-surfaces in $\mathcal{X}_{0}$. By Equations (6.10)-(6.11), we have $\vartheta_{p}|_{D_{t}}=0$ if $p\notin\partial P_{\mathbb{Z}}$. On the other hand, by Lemma 6.35, $D_{t}$ is a cycle of rational curves, with 0-dimensional strata $x_{v}$ in one-to-one correspondence with the integral points $v$ of $\partial P$ corresponding to the 0-surfaces in $\mathcal{X}_{0}$. Moreover, the theta functions $\vartheta_{p}$ with $p\in\partial P$ restrict to toric monomials on the irreducible components of $D_{t}$. It follows that for every 0-dimensional stratum $x_{v}$ of $D_{t}$, we have $\vartheta_{p}(x_{v})\neq 0$ if $p=v$, and $\vartheta_{p}(x_{v})=0$ if $p\in P_{\mathbb{Z}}\setminus\{v\}$. Hence, we have $\sum_{p\in P_{v}}\vartheta_{p}(x_{v})\neq 0$, and so $x_{v}\notin C_{t}$. This concludes the proof of i).